On boundary value problems for one-dimensional Schrödinger operator

Alexander Makin Moscow State Academy of Instrument-Making and Informatics

We will consider nonselfadjoint boundary value problems for one-dimensional Schrodinger operator u′′ − q(x)u defined on the segment [0, 1]. We establish sufficient conditions of summability of spectral expansions which hold for many types of boundary conditions, for instance, Birkhoff-regular but not strongly regular, irregular, multipoint boundary conditions etc. A lot of interesting examples will be considered.